OpenStudy (anonymous):

use substitution to solve the system {2x-5y=-5, -2x+8y=-58

OpenStudy (amistre64):

id rather just use common sense

OpenStudy (amistre64):

3y = 63; y = 21 x = 10

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

actually it looks like you get 3y = -63 so y = -21 yes?

OpenStudy (anonymous):

\[2x-5y=-5\] \[-2x+8y=-58\] add to get \[3y=-63\] \[y=-21\] substitute back to get \[2x-5\times -21=-5\] \[2x+105=-5\] \[2x=-110\] \[x=-55\]

OpenStudy (amistre64):

lol.... did i drop a negative